Please DO NOT use excel. Show all steps please.

You have 2 mutually exclusive alternatives and a MARR of 9%. Which alternative is preferred, based on repeatability assumption?

Alternative E F
Capital Investment $14,000 $65,000
Annual Expenses $14,000 $9,000
Useful Life (years) 4 20
Market Value at end of useful life $8,000 $13,000
Answer #1

MARR = 9%

EUAC of E = 14000*(A/P,9%,4) + 14000 - 8000*(A/F,9%,4)

= 14000*0.308669 + 14000 - 8000*0.218669

= 16572.01

EUAC of F = 65000*(A/P,9%,20) + 9000 - 13000*(A/F,9%,20)

= 65000*0.109546 + 9000 - 13000*0.019546

= 15866.42

As EUAC of F is lower, it should be selected